OpenStudy (anonymous):

Suppose A is n*n square matrice on R that A^2=-I proof : 1. A has no Eigenvalues on R . 2. n is even

OpenStudy (anonymous):

hey friend, A is a matrix how A^2 =-1? is there something missed?

OpenStudy (anonymous):

I am studying this part, too. so that's why I concern. We can do the stuff together.

OpenStudy (anonymous):

that is not -1 is -I (identity matrice)

OpenStudy (anonymous):

I mean A^2+I=0

OpenStudy (anonymous):

oh, ok, I misunderstand, hihi, sorry, I'm thinking.

OpenStudy (anonymous):

hi friend, I think we can start at A^2 +I = 0, since I is identity matrix, so A^2 must have the form of -1 in main diagonal line and any other entries =0. and then A^2 = A *A it's mean -1= something^2 it's mean...hehehe...what? you know and I know , too

OpenStudy (anonymous):

A^2 =\[\left[\begin{matrix}-1 & 0 &0 &..... \\ 0 & -1 & 0 & ..........\\0 & 0 & -1 & .....\end{matrix}\right]\]

OpenStudy (anonymous):

do you get what I mean? if my English is toooo complicated, and you don't understand, just speak out. I'm ok

OpenStudy (anonymous):

I think do not understand

OpenStudy (anonymous):

@Amirsd , you there? sorry, my computer lost connection with the site a while

OpenStudy (anonymous):

Yes I am here friend

OpenStudy (anonymous):

ok, I explain you now, you have A^2 + I = 0 right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\left[\begin{matrix}-1 & 0 &0&......\\ 0 & -1&0&.........\\0&0&-1&......\end{matrix}]+\left[\begin{matrix}1 & 0&0&.....\\ 0 & 1&0&...\\0&0&1&...\end{matrix}\right]\right]\] =0

OpenStudy (anonymous):

the first matrix is A^2

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

the second one is I , sum of them is 0

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

so, A^2 = A * A what it mean? it mean A is a complex matrix with entries in main diagonal line is . Why? because i^2 = -1

OpenStudy (anonymous):

entries = i

OpenStudy (anonymous):

yes that is right

OpenStudy (anonymous):

ok, can you make up the characteristic equation of \[\left[\begin{matrix}i & 0&0&.... \\ 0& i&0&...\\0&0&i&.....\end{matrix}\right]\]

OpenStudy (anonymous):

and get a real eigenvalue?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

that the proof of your a) question. no real eigenvalue for A

OpenStudy (anonymous):

so what about 2 ?

OpenStudy (anonymous):

you mean question b) n must be even?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

because when you diagonalize a matrix to get A^n and L^n (lamda) to the theorem, you have L^n is eigenvalue of A^n , since -1 =i^2 , that means n =2 if n =3 i^3 is not -1, it's is 1 ---> A^2 + I != 0 (not equal )

OpenStudy (anonymous):

got it? hehehe... thanks for your question, I gain too much

OpenStudy (anonymous):

so \[AX=\lambda A\] why whe diagonalize ?

OpenStudy (anonymous):

I mean for get A^n?

OpenStudy (anonymous):

our matrice was A^2 is't it ?

OpenStudy (anonymous):

hey, friend, your problem relate to that, A^2 if you don't diagonalize how to come up with A^2 by D and no way to deal with Lamda.

OpenStudy (anonymous):

so what about A^n why you write that ?

OpenStudy (anonymous):

you see your A is a diagonal matrix, it mean A^2 = A *A = {diagonal entries ^2} and A^n = {diagonal entries ^n} and the eigenvalue is counted by L ^2, L ^n .... respectively,. If n= even , i ^ (even) number = -1, if n = odd, i^(odd) =1 which not satisfy the condition of A^n + I =0

OpenStudy (anonymous):

I have to go to my friend' post to help her.

OpenStudy (anonymous):

thank you for helping friend I took your time

OpenStudy (anonymous):

it's ok, I help you to learn too.